OpenStudy (anonymous):

Use inductive reasoning to predict the next number in the sequence. 3, 10, 17, 24, 31 A. 37 B. 34 C. 38 D. 45

OpenStudy (jadeishere):

38 it goes up by 7 each time 10 - 3 = 7 17 - 10 = 7 24 - 17 = 7 etc.

OpenStudy (anonymous):

so A right

OpenStudy (anonymous):

or is it C

OpenStudy (jadeishere):

C

OpenStudy (anonymous):

The numbers are increasing by 7, right?

OpenStudy (anonymous):

thanks I didn't see ur answer clearly smh 0_o

OpenStudy (anonymous):

The last number shown in the sequence is 31, so you add 7 to that which is 38.

OpenStudy (jadeishere):

Yep